OpenStudy (webo):

Juliana has created the function f(x) = the quantity of 3x plus 2, divided by 4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that, using the inverse function to solve for x = 24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences.

jimthompson5910 (jim_thompson5910):

The function is this right? \[\Large f(x) = \frac{3x+2}{4}\]

OpenStudy (webo):

I can explain the answer just help me how to find out f(x) = 3x + 2/ 4

OpenStudy (legomyego180):

Those would be really expensive text messages.

OpenStudy (webo):

@jim_thompson5910 yes

jimthompson5910 (jim_thompson5910):

Are you familiar with how to find the inverse of a function?

OpenStudy (webo):

Somewhat

jimthompson5910 (jim_thompson5910):

The basic outline is to step 1) replace f(x) with y step 2) swap x and y step 3) solve for y

OpenStudy (webo):

I'm stuck on solve for y part

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{3x+2}{4}\] \[\Large y = \frac{3x+2}{4}\] \[\Large x = \frac{3y+2}{4}\] \[\Large 4x = 4*\frac{3y+2}{4}\] \[\Large 4x = 3y+2\] Does this help you see how to solve for y?

OpenStudy (webo):

4x - 2 = 3y?

jimthompson5910 (jim_thompson5910):

yes, now what comes next?

OpenStudy (webo):

I'm not sure...

jimthompson5910 (jim_thompson5910):

isolate the y by undoing the multiplication of 3

jimthompson5910 (jim_thompson5910):

3y means 3 times y

jimthompson5910 (jim_thompson5910):

what operation is the opposite of multiplication?

OpenStudy (webo):

Division, I'm trying to solve it on a paper.

jimthompson5910 (jim_thompson5910):

yes, so divide both sides by 3

OpenStudy (webo):

4x -2 / 3 = 3y / 3?

jimthompson5910 (jim_thompson5910):

\[\Large 4x-2 = 3y\] \[\Large \frac{4x-2}{3} = \frac{3y}{3}\] \[\Large \frac{4x-2}{3} = y\] \[\Large y = \frac{4x-2}{3}\] So the inverse function is \[\Large f^{-1}(x) = \frac{4x-2}{3}\]

jimthompson5910 (jim_thompson5910):

once you get the inverse function, plug in x = 24. The output will be the number of texts you can do (if you only have $24 to spend)

OpenStudy (webo):

f-1(24) = 4(24) - 2 / 3?

jimthompson5910 (jim_thompson5910):

yes, but make sure to use parentheis \[\Large f^{-1}(x) = \frac{4x-2}{3}\] \[\Large f^{-1}(24) = \frac{4*24-2}{3}\] \[\Large f^{-1}(24) = ???\]

jimthompson5910 (jim_thompson5910):

use parenthesis instead of writing ` 4(24) - 2 / 3` write ` (4(24) - 2) / 3` instead to indicate that the "4(24)-3" portion is all over the 3

OpenStudy (webo):

\[f ^{-1}(24) = \frac{ 96 }{ 3 }\]

jimthompson5910 (jim_thompson5910):

you forgot about the "-2" up top

OpenStudy (webo):

\[f ^{-1}(24)=\frac{ 94 }{ 3 }\]

jimthompson5910 (jim_thompson5910):

what is 94/3 as a decimal?

OpenStudy (webo):

31.3333333

jimthompson5910 (jim_thompson5910):

that will round down to 31 so the most text messages you can send is 31. If you send 32, then you'll go over $24

OpenStudy (webo):

That's the answer?

jimthompson5910 (jim_thompson5910):

Well it looks like the teacher wants you to explain the whole process. Your teacher doesn't seem to be concerned with one simple numeric answer.

OpenStudy (webo):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome